# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ 3.2. Lagrangian chaos for stochastic PEs
Let $Q$ be as in Subsection 2.2.2. Then there exists a deterministic constant $\lambda_{+}>0$ such that, for every initial position $\mathbf{x}\in\mathbb{T}^{3}$ and initial velocity $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$,
@@
Moreover, the flow expands exponentially in all directions with rate $\lambda$, i.e., for every $\mathbf{x}\in\mathbb{T}^{3}$, $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$ and $\xi\in{\mathbb{R}}^{3}\setminus\{0\}$,
@@
(3.6) $&\partial_{t}f_{t}+(u_{t}\cdot\nabla f_{t})=0&\text{ on}&\mathbb{T}^{3},\\ &f_{0}=f&\text{ on }&\mathbb{T}^{3},\end{aligned}\right.$
where $f$ is given. Due to (3.1) and $\sigma>\frac{9}{2}$, we have $(\phi_{t})_{t}\in C({\mathbb{R}}_{+};C^{2}(\mathbb{T}^{3}))$ a.s. and therefore the unique regular (in the Sobolev sense) solution to (3.6) is given by $f_{t}(\mathbf{x})=f(\phi_{t}^{-1}(\mathbf{x}))$. Since $\phi_{t}$ is volume preserving, it follows that
(3.7) $\|\nabla f_{t}\|_{L^{1}(\mathbb{T}^{3})}=\int_{\mathbb{T}^{3}}|(\nabla\phi_{t}(\mathbf{x}))^{-\top}\nabla f(\mathbf{x})|\,\mathrm{d}\mathbf{x}.$
The RHS of the above suggests investigating the properties of $(\nabla\phi_{t}(\mathbf{x}))^{-\top}$. In analogy with Theorem (3.5), we have
@@
Let $Q$ be as in Subsection 2.2.2. Let $\overline{\lambda}_{+}>0$ be as in Theorem 3.6, and fix $p\in[1,\infty)$. Then, for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$, $\eta\in(0,\overline{\lambda}_{+})$ and a non-zero $f\in W^{1,p}(\mathbb{T}^{3})$ such that $\int_{\mathbb{T}^{3}}f(\mathbf{x})\,\mathrm{d}\mathbf{x}=0$, there exists a constant $r=r(v,f,\eta,p)>0$ such that
(3.8) $\|\nabla f_{t}\|_{L^{p}(\Omega\times\mathbb{T}^{3})}\geq r\,e^{\eta t}\ \text{for all }t>0.$
@@
It suffices to prove (3.8) for $p=1$. Let $\mathcal{O}_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{\mathbf{x}\in\mathbb{T}^{3}\,:\,\nabla f(\mathbf{x})\neq 0\}$. By assumption, $\mathrm{Leb}(\mathcal{O}_{0})>0$. Theorem 3.6 ensures that, for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$ and $\mathbf{x}\in\mathcal{O}$,
@@
where $\overline{\lambda}_{+}>0$ is deterministic. For all $\eta\in(0,\overline{\lambda}_{+})$, $\omega\in\Omega$ and $\mathbf{x}\in\mathbb{T}^{3}$, set
$R(\omega,\mathbf{x})\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathbf{1}_{\mathcal{O}_{0}}(\mathbf{x})\,\textstyle{\inf}_{t>0}\big{(}|\nabla\phi^{t}_{\omega,v}(\mathbf{x})\nabla f(\mathbf{x})|e^{-\eta t}\big{)}.$
It follows from (3.9) that $\mathbf{P}\times\mathrm{Leb}(R>0)=\mathrm{Leb}(\mathcal{O}_{0})>0.$ Due to (3.7) and the above reasoning, the estimate (3.8) for $p=1$ holds with $r=\mathbf{E}\int_{\mathbb{T}^{3}}R(\cdot,\mathbf{x})\,\mathrm{d}\mathbf{x}>0$. ∎
@@
The proofs of Theorems 3.5 and 3.6 are given in Subsection 3.3 based on sufficient conditions for positivity of the top Lyapunov exponents proven in [16, Section 4] and further results on the stochastic PEs (1.1), see Subsection 3.3 below. Sections 4-7 are devoted to the proofs of the above-mentioned additional results on (1.1).
@@
As mentioned in Section 1, the proof of Theorems 3.5 and 3.6 are based on the sufficient condition for the positivity of the top Lyapunov exponent proven in [16, Section 4]. For brevity, here we limit ourselves to explaining how to apply their results to the current situation rather than presenting the findings in [16] in full generality. In this subsection, we will state several results which will be proven in the later sections. Combining the latter with the results in [16], we will prove Theorems 3.5 and 3.6. We begin by focusing on checking the assumptions in [16, Proposition 4.17]. Let us begin by introducing some notation and terminology.
@@
A mapping $\mathcal{R}:[0,\infty)\times\Omega\times Z\to Z$ on a polish space $Z$, written as $(t,\omega,z)\mapsto\mathcal{R}_{\omega}^{t}(z)$, is said to be a random dynamical system (RDS in the following) if the following are satisfied (recall that $\mathscr{B}$ denotes the Borel $\sigma$-algebra):
@@
  • (Uniform continuity) For all $\omega\in\Omega$ and all bounded set $U\subseteq Z$, the set $\mathcal{R}_{\omega}(U)\subseteq Z$ and $\mathcal{R}_{\omega}(U)$ is bounded in $Z$.
@@
The above is a special case of Proposition 5.4 whose proof is postponed to Subsection 5.2. We now return to the proof of Theorems 3.5 and 3.6. We begin with the proof of the former, as the latter can be derived by replacing $\nabla\phi_{t}(\mathbf{x})$ with $(\nabla\phi_{t}(\mathbf{x}))^{-\top}$ in the arguments presented below.
To prove Theorem 3.5 we would like to apply [16, Propositions 3.16 and 4.17] on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$, with the above choice of $(\theta^{t})_{t}$ and
@@
where $\omega\in\Omega$, $t\geq 0$, $\mathbf{x}\in\mathbb{T}^{3}$ and $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$. Moreover, for $\omega\in\Omega$, $(\phi^{t}_{\omega})_{t}$ denotes the Lagrangian flow associated with the RDS in Proposition 3.8:
(3.13) $\partial_{t}\phi^{t}_{\omega}(\mathbf{x})=\mathcal{U}^{t}_{\omega}(\phi^{t}_{\omega}(\mathbf{x}))\ \text{ for }t\geq 0,\qquad\phi^{0}_{\omega}(\mathbf{x})=\mathbf{x}\in\mathbb{T}^{3},$
@@
and $w(\cdot)$ is as in (2.1). As commented below (2.14), the matrix $\mathcal{A}_{\omega,(v,\mathbf{x})}$ is well defined. Moreover, the fact that $\nabla\cdot\mathcal{U}^{t}_{\omega}=0$ implies $\mathcal{A}_{\omega,(v,\mathbf{x})}^{t}\in\mathrm{SL}_{3}({\mathbb{R}})$ (i.e., $\mathcal{A}_{\omega,(v,\mathbf{x})}$ is invertible) for a.a. $\omega\in\Omega$ and all $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$.
Let us emphasize that, for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$, we have $\mathcal{V}^{t}_{\omega}(v)=v_{t}(\omega)$ for a.a. $\omega\in\Omega$ and $t\geq 0$. Hence, the above definition of $\phi^{t}_{\omega}$ coincides with the one in Theorem 1.1 whenever the $\Omega$-expectional sets are allowed to depend on the initial data $v$ (as in the case of the statement in Theorem 1.1).
Proposition 3.8 implies that $\mathcal{T}$ is a random dynamical system on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$ as well, and $\mathcal{A}$ is a (three-dimensional) linear cocycle over $\mathcal{T}$ ([16, Definition 3.11]). Moreover, Proposition 3.8 also ensures that the assumptions [16, (H1) and (H3)] are satisfied. Next, we verify the integrability conditions [16, (H2)] for the linear cocycle $\mathcal{A}$, i.e., (here $\log^{+}\cdot=\max\{\log\cdot,0\}$)
(3.15) $\mathbf{E}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathbb{T}^{3}}\sup_{0\leq t\leq 1}\log^{+}|\mathcal{A}_{(v,\mathbf{x})}^{t}|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)$ <<$\infty$,< ∞ ,
(3.16) $\mathbf{E}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathbb{T}^{3}}\sup_{0\leq t\leq 1}\log^{+}|(\mathcal{A}_{(v,\mathbf{x})}^{t})^{-1}|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)$ <<$\infty$.< ∞ .
@@
(3.17) $\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathbb{T}^{3}}|\nabla u(\mathbf{x})|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)\leq C(\mathcal{E}),$
@@
Next, we show how Proposition 3.9 implies (3.15)-(3.16) with $\mathcal{A}_{\omega,(v,\mathbf{x})}^{t}$ as in (3.12). Differentiating (3.13), we obtain
(3.18) $\partial_{t}\mathcal{A}_{\omega,(v,\mathbf{x})}^{t}=\nabla\mathcal{U}^{t}_{\omega}(\phi^{t}_{\omega}(\mathbf{x}))\,\mathcal{A}_{\omega,(v,\mathbf{x})}^{t}\ \text{ for }t\geq 0,\qquad\mathcal{A}_{\omega,(v,\mathbf{x})}^{0}=\mathrm{Id}.$
@@
$\mathbf{E}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathbb{T}^{3}}\int_{0}^{1}|\nabla u_{t}(\mathbf{x}_{t})|\,\mathrm{d}t\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}$ =$\stackrel{{\scriptstyle(i)}}{{=}}\mathbf{E}\int_{0}^{1}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathbb{T}^{3}}|\nabla u_{t}(\mathbf{x})|\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}\,\,\mathrm{d}t$
@@
where in $(i)$ and $(ii)$ we used that transformation $\mathbf{x}\mapsto\mathbf{x}_{t}=\phi_{t}(\mathbf{x})$ is measure preserving and the invariance of $\mu$, respectively. Finally, (3.16) follows similarly as
@@
In the following, if no confusion seems likely, we will not distinguish between $(v_{t})_{t}$ and the generated RDS $(\mathcal{V}^{t}(v))_{t}$, and similarly for $(u_{t})_{t}$ and $(\mathcal{U}^{t}(v))_{t}$.
@@
with initial data $\xi_{0}=\xi\in\mathbb{S}^{2}$ and $\widecheck{\xi}_{0}=\widecheck{\xi}\in\mathbb{S}^{2}$; respectively. Here, $\mathbb{S}^{2}\subseteq{\mathbb{R}}^{3}$ is the two-dimensional sphere, $\Pi_{\xi}=\mathrm{Id}-(\xi\otimes\xi)$ is the orthogonal projection from ${\mathbb{R}}^{3}$ onto the tangent space of $\mathbb{S}^{2}$ at the point $\xi$, and $\top$ denotes the transposition.
With a slight abuse of notation, we still denote by $\xi_{t}$ (resp., $\widecheck{\xi}_{t}$) the equivalent class in the two-dimensional projective space $\mathrm{Pr}_{2}$ given by $\pi(\xi_{t})$ (resp., $\pi(\widecheck{\xi}_{t})$) where $\pi:\mathbb{S}^{2}\to\mathrm{Pr}_{2}$ is the usual quotient map which identifies opposite points on $\mathbb{S}^{2}$.
The remaining ingredients needed to apply [16, Proposition 4.17] are provided by the following three results. In the following, for a metric space $Z$, we denote by $B_{Z}(z,\varepsilon)$ the ball of radius $\varepsilon>0$ with center $z\in Z$. The superscript ${\rm c}$ indicates the complement of a set. We denote by $L_{t}((v,\mathbf{x}),\cdot)$ the Markov transition function of the Lagrangian process $(v_{t},\mathbf{x}_{t})_{t}$, i.e., for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$, $\mathbf{x}\in\mathbb{T}^{3}$ and $t>0$,
@@
Let Assumption 2.2 be satisfied and let $\sigma$ be as in (2.14). The following hold for all $t>0$ and $\mathbf{a}\in\mathcal{O}$.
@@
    The support of $L_{t}((0,\mathbf{a}),\cdot)$ contains $\{0\}\times\mathbb{T}^{3}$.
@@
    $\mathbf{P}\big{(}(v_{t},\mathbf{x}_{t},\xi_{t})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathbb{T}^{3}}(\mathbf{a},\varepsilon)\times B_{\mathrm{Pr}_{2}}(\eta^{\prime},\varepsilon)\,|\,(v_{0},\mathbf{x}_{0},\xi_{0})=(0,\mathbf{a},\eta)\big{)}$ $>0,$
    $\mathbf{P}\big{(}(v_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathbb{T}^{3}}(\mathbf{a},\varepsilon)\times B_{\mathrm{Pr}_{2}}(\eta^{\prime},\varepsilon)\,|\,(v_{0},\mathbf{x}_{0},\widecheck{\xi}_{0})=(0,\mathbf{a},\eta)\big{)}$ $>0.$
@@
    $\mathbf{P}\big{(}(v_{t},\mathbf{x}_{t},\mathcal{A}_{t})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathbb{T}^{3}}(\mathbf{a},\varepsilon)\times B_{{\mathbb{R}}^{3\times 3}}^{{\rm c}}(0,M)\,|\,(v_{0},\mathbf{x}_{0},\mathcal{A}_{0})=(0,\mathbf{a},\mathrm{Id})\big{)}>0.$
@@
Let Assumption 2.2 be satisfied and let $\sigma$ be as in (2.14). Then the Markov semigroups associated with the processes $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ and $(v_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})_{t}$ are strong Feller in $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}\times\mathrm{Pr}_{2}.$
Proposition 3.12 (Weak irreducibility – Projective processes).
@@
  • The support of any stationary measure for the Lagrangian process $(u_{t},\mathbf{x}_{t})$ on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$ must contain $\{0\}\times\mathbb{T}^{3}$.
  • The support of any stationary measure for the projective processes $(u_{t},\mathbf{x}_{t},\xi_{t})$ and $(u_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})$ on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}\times\mathrm{Pr}_{2}$ must contain $\{0\}\times\mathbb{T}^{3}\times\mathrm{Pr}_{2}$.
Propositions 3.11 and 3.12 readily yield the following result.
Corollary 3.13 (Uniqueness of invariant measures for auxiliary processes).

### v2

@@ 3.2. Lagrangian chaos for stochastic PEs
Let $Q$ be as in Subsection 2.2.2. Then there exists a deterministic constant $\lambda_{+}>0$ such that, for every initial position $\mathbf{x}\in\mathcal{O}$ and initial velocity $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$,
@@
Moreover, the flow expands exponentially in all directions with rate $\lambda$, i.e., for every $\mathbf{x}\in\mathcal{O}$, $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$ and $\xi\in{\mathbb{R}}^{3}\setminus\{0\}$,
@@
(3.6) $&\partial_{t}f_{t}+(u_{t}\cdot\nabla f_{t})=0&\text{ on}&\mathcal{O},\\ &f_{0}=f&\text{ on }&\mathcal{O},\end{aligned}\right.$
where $f$ is given. Due to (3.1) and $\sigma>\frac{9}{2}$, we have $(\phi_{t})_{t}\in C({\mathbb{R}}_{+};C^{2}(\overline{\mathcal{O}}))$ a.s. and therefore the unique regular (in the Sobolev sense) solution to (3.6) is given by $f_{t}(\mathbf{x})=f(\phi_{t}^{-1}(\mathbf{x}))$. Since $\phi_{t}$ is volume preserving, it follows that
(3.7) $\|\nabla f_{t}\|_{L^{1}(\mathcal{O})}=\int_{\mathcal{O}}|(\nabla\phi_{t}(\mathbf{x}))^{-\top}\nabla f(\mathbf{x})|\,\mathrm{d}\mathbf{x}.$
The RHS of the above suggests investigating the properties of $(\nabla\phi_{t}(\mathbf{x}))^{-\top}$ (here, $\top$ denotes the transpose). In analogy with Theorem (3.5), we have
@@
Let $Q$ and $\overline{\lambda}_{+}>0$ be as in Subsection 2.2.2 and Theorem 3.6, respectively. Then, for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$, $\eta\in(0,\overline{\lambda}_{+})$ and a non-zero $f\in W^{1,1}(\mathcal{O})$ such that $\int_{\mathcal{O}}f(\mathbf{x})\,\mathrm{d}\mathbf{x}=0$, there exists a constant $R=R(v,f,\eta)>0$ such that, for all $t>0$,
(3.8) $\mathbf{E}\int_{\mathcal{O}}f_{t}(\mathbf{x})\,\mathrm{d}\mathbf{x}\geq R\,e^{\eta t}.$
@@
Let $\mathcal{O}_{f}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{\mathbf{x}\in\mathcal{O}\,:\,\nabla f(\mathbf{x})\neq 0\}$. By assumption, $\mathrm{Leb}(\mathcal{O}_{f})>0$. Theorem 3.6 ensures that, for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$ and $\mathbf{x}\in\mathcal{O}$,
@@
where $\overline{\lambda}_{+}>0$ is deterministic. For all $\eta\in(0,\overline{\lambda}_{+})$, $\omega\in\Omega$ and $\mathbf{x}\in\mathcal{O}$, set
$r(\omega,\mathbf{x})\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathbf{1}_{\mathcal{O}_{f}}(\mathbf{x})\,\textstyle{\inf}_{t>0}\big{(}|\nabla\phi^{t}_{\omega,v}(\mathbf{x})\nabla f(\mathbf{x})|e^{-\eta t}\big{)}.$
It follows from (3.9) that $\mathbf{P}\times\mathrm{Leb}(r>0)=\mathrm{Leb}(\mathcal{O}_{f})>0.$ Due to (3.7) and the above reasoning, the estimate (3.8) holds with $R=\mathbf{E}\int_{\mathcal{O}}r(\mathbf{x})\,\mathrm{d}\mathbf{x}>0$. ∎
@@
The proofs of Theorems 3.5 and 3.6 are given in Subsection 3.3 based on sufficient conditions for positivity of the top Lyapunov exponents proven in [16, Section 4] and further results on the stochastic PEs (1.1). Details are given in Subsection 3.3 below. Sections 4-7 are devoted to the proofs of the above-mentioned additional results on (1.1).
@@
As mentioned in Section 1, the proofs of Theorems 3.5 and 3.6 are based on the sufficient condition for the positivity of the top Lyapunov exponent proven in [16, Section 4]. For brevity, here we limit ourselves to explaining how to apply their framework to the current situation rather than presenting the results in [16] in full generality. In this subsection, we will state several results which will be proven in the later sections. Combining the latter with the results in [16], we will prove Theorems 3.5 and 3.6. We begin by focusing on checking the assumptions in [16, Proposition 4.17]. Let us begin by introducing some notation and terminology.
@@
A mapping $\mathcal{R}:[0,\infty)\times\Omega\times Z\to Z$ on a polish space $Z$, written as $(t,\omega,z)\mapsto\mathcal{R}_{\omega}^{t}(z)$, is said to be a (continuous) random dynamical system (RDS in the following) if the following are satisfied (recall that $\mathscr{B}$ denotes the Borel $\sigma$-algebra):
@@
  • (Boundedness and uniform continuity) For all $\omega\in\Omega$ and all bounded set $U\subseteq[0,\infty)\times Z$, the set $\mathcal{R}_{\omega}(U)$ is bounded in $Z$ and uniformly continuous.
@@
The above is a special case of Proposition 5.4 whose proof is given in Subsection 5.2. We now return to the proof of Theorems 3.5 and 3.6. We begin with the proof of the former, as the latter can be derived by replacing $\nabla\phi_{t}(\mathbf{x})$ with $(\nabla\phi_{t}(\mathbf{x}))^{-\top}$ in the arguments presented below.
To prove Theorem 3.5, we apply [16, Propositions 3.16 and 4.17] on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$, with the above choice of $(\theta^{t})_{t}$ and
@@
where $\omega\in\Omega$, $t\geq 0$, $\mathbf{x}\in\mathcal{O}$ and $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$. Moreover, for $\omega\in\Omega$, $(\phi^{t}_{\omega})_{t}$ denotes the Lagrangian flow associated with the RDS in Proposition 3.8:
(3.13) $\partial_{t}\phi^{t}_{\omega}(\mathbf{x})=\mathcal{U}^{t}_{\omega}(\phi^{t}_{\omega}(\mathbf{x}))\ \text{ for }t\geq 0,\qquad\phi^{0}_{\omega}(\mathbf{x})=\mathbf{x}\in\mathcal{O},$
@@
and $w(\cdot)$ is as in (2.1). As commented below (2.14), the matrix $\mathcal{A}_{\omega,(v,\mathbf{x})}$ is well defined. Moreover, the fact that $\nabla\cdot\mathcal{U}^{t}_{\omega}=0$ implies $\mathcal{A}_{\omega,(v,\mathbf{x})}^{t}\in\mathrm{SL}_{3}({\mathbb{R}})$ (i.e., $\mathcal{A}_{\omega,(v,\mathbf{x})}$ is invertible) for a.a. $\omega\in\Omega$ and all $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$.
Let us recall that, for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$, we have $\mathcal{V}^{t}_{\omega}(v)=v_{t}(\omega)$ for a.a. $\omega\in\Omega$ and $t\geq 0$. Hence, the above definition of $\phi^{t}_{\omega}$ coincides with the one in Theorem 1.1 whenever the $\Omega$-expectional sets are allowed to depend on the initial data $v$ (as in the case of the statement in Theorem 1.1). In the following, if no confusion seems likely, we will not distinguish between $(v_{t})_{t}$ and the generated RDS $(\mathcal{V}^{t}(v))_{t}$, and similarly for $(u_{t})_{t}$ and $(\mathcal{U}^{t}(v))_{t}$.
Proposition 3.8 implies that $\mathcal{T}$ is a random dynamical system on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$ as well, and $\mathcal{A}$ is a (three-dimensional) linear cocycle over $\mathcal{T}$ ([16, Definition 3.11]). Moreover, Proposition 3.8 also ensures that the assumptions [16, (H1) and (H3)] are satisfied. Next, we verify the integrability conditions [16, (H2)] for the linear cocycle $\mathcal{A}$, i.e., (here $\log^{+}\cdot=\max\{\log\cdot,0\}$)
(3.15) $\mathbf{E}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathcal{O}}\sup_{0\leq t\leq 1}\log^{+}|\mathcal{A}_{(v,\mathbf{x})}^{t}|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)$ <<$\infty$,< ∞ ,
(3.16) $\mathbf{E}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathcal{O}}\sup_{0\leq t\leq 1}\log^{+}|(\mathcal{A}_{(v,\mathbf{x})}^{t})^{-1}|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)$ <<$\infty$.< ∞ .
@@
(3.17) $\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathcal{O}}|\nabla u(\mathbf{x})|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)\leq C(\mathcal{E}),$
@@
Next, we show how Proposition 3.9 implies (3.15)-(3.16) with $\mathcal{A}_{(v,\mathbf{x})}^{t}=\mathcal{A}_{\cdot,(v,\mathbf{x})}^{t}$ as in (3.12). Differentiating (3.13), we obtain
(3.18) $\partial_{t}\mathcal{A}_{(v,\mathbf{x})}^{t}=\nabla u_{t}(\phi^{t}(\mathbf{x}))\,\mathcal{A}_{(v,\mathbf{x})}^{t}\ \text{ for }t\geq 0,\qquad\mathcal{A}_{(v,\mathbf{x})}^{0}=\mathrm{Id}.$
@@
$\mathbf{E}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathcal{O}}\int_{0}^{1}|\nabla u_{t}(\mathbf{x}_{t})|\,\mathrm{d}t\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}$ =$\stackrel{{\scriptstyle(i)}}{{=}}\mathbf{E}\int_{0}^{1}\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})}\int_{\mathcal{O}}|\nabla u_{t}(\mathbf{x})|\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}\,\,\mathrm{d}t$
@@
where in $(i)$ and $(ii)$ we used that transformation $\mathbf{x}\mapsto\mathbf{x}_{t}=\phi_{t}(\mathbf{x})$ is volume preserving and the invariance of $\mu$, respectively. Finally, (3.16) follows similarly as
@@
with initial data $\xi_{0}=\xi\in\mathbb{S}^{2}$ and $\widecheck{\xi}_{0}=\widecheck{\xi}\in\mathbb{S}^{2}$; respectively. Here, $\mathbb{S}^{2}\subseteq{\mathbb{R}}^{3}$ is the two-dimensional sphere and $\Pi_{\xi}=\mathrm{Id}-(\xi\otimes\xi)$ is the orthogonal projection from ${\mathbb{R}}^{3}$ onto the tangent space of $\mathbb{S}^{2}$ at the point $\xi$. With a slight abuse of notation, we still denote by $\xi_{t}$ (resp., $\widecheck{\xi}_{t}$) the equivalent class in the two-dimensional projective space $\mathrm{Pr}_{2}$ given by $\pi(\xi_{t})$ (resp., $\pi(\widecheck{\xi}_{t})$) where $\pi:\mathbb{S}^{2}\to\mathrm{Pr}_{2}$ is the usual quotient map which identifies opposite points on $\mathbb{S}^{2}$.
The remaining ingredients needed to apply [16, Proposition 4.17] are provided by the following three results. In the following, for a metric space $Z$, we denote by $B_{Z}(z,\varepsilon)$ the ball of radius $\varepsilon>0$ with center $z\in Z$. The superscript ${\rm c}$ indicates the complement of a set. We denote by $L_{t}((v,\mathbf{x}),\cdot)$ the Markov transition function of the Lagrangian process $(v_{t},\mathbf{x}_{t})_{t}$, i.e., for all $v\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$, $\mathbf{x}\in\mathcal{O}$ and $t>0$,
@@
Let Assumption 2.2 be satisfied and let $\sigma$ be as in (2.14). The following results hold for all $t>0$ and $\mathbf{a}\in\mathcal{O}$.
@@
    The support of $L_{t}((0,\mathbf{a}),\cdot)$ contains $\{0\}\times\mathcal{O}$.
@@
    $\mathbf{P}\big{(}(v_{t},\mathbf{x}_{t},\xi_{t})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathcal{O}}(\mathbf{a},\varepsilon)\times B_{\mathrm{Pr}_{2}}(\eta^{\prime},\varepsilon)\,|\,(v_{0},\mathbf{x}_{0},\xi_{0})=(0,\mathbf{a},\eta)\big{)}$ $>0,$
    $\mathbf{P}\big{(}(v_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathcal{O}}(\mathbf{a},\varepsilon)\times B_{\mathrm{Pr}_{2}}(\eta^{\prime},\varepsilon)\,|\,(v_{0},\mathbf{x}_{0},\widecheck{\xi}_{0})=(0,\mathbf{a},\eta)\big{)}$ $>0.$
@@
    $\mathbf{P}\big{(}(v_{t},\mathbf{x}_{t},\mathcal{A}_{t})\in B_{\mathbb{H}^{\sigma}}(0,\varepsilon)\times B_{\mathcal{O}}(\mathbf{a},\varepsilon)\times B_{{\mathbb{R}}^{3\times 3}}^{{\rm c}}(0,M)\,|\,(v_{0},\mathbf{x}_{0},\mathcal{A}_{0})=(0,\mathbf{a},\mathrm{Id})\big{)}>0.$
@@
Let Assumption 2.2 be satisfied and let $\sigma$ be as in (2.14). Then the Markov semigroups associated with the processes $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ and $(v_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})_{t}$ are strong Feller in $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}\times\mathrm{Pr}_{2}.$
Proposition 3.12 (Weak irreducibility – Auxiliary processes).
@@
  • The support of any stationary measure for the Lagrangian process $(u_{t},\mathbf{x}_{t})$ on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$ must contain $\{0\}\times\mathcal{O}$.
  • The support of any stationary measure for the projective processes $(u_{t},\mathbf{x}_{t},\xi_{t})$ and $(u_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})$ on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}\times\mathrm{Pr}_{2}$ must contain $\{0\}\times\mathcal{O}\times\mathrm{Pr}_{2}$.
Propositions 3.11, 3.12, and results from Appendix A yield the following uniqueness result for ergodic measures of the Lagrangian and projective processes.
Corollary 3.13 (Uniqueness of ergodic measures – Auxiliary processes).